### v1

# From Acceleration to Saturation: Scaling Behavior of Bootstrapped Language Model Pretraining

Abstract
Bootstrapped pretraining, i.e., the reuse of a pretrained base model for further pretraining, such as continual pretraining or model growth, is promising at reducing the cost of training language models from scratch. However, its effectiveness remains unclear, especially when applied to overtrained base models. In this work, we empirically study the scaling behavior of bootstrapped pretraining and find that its scaling efficiency diminishes in a predictable manner: The scaling exponent with respect to second-stage pretraining tokens decreases logarithmically with the number of tokens used to pretrain the base model. The joint dependence on first- and second-stage tokens is accurately modeled by a simple scaling law. Such saturation effect reveals a fundamental trade-off in multi-stage pretraining strategies: the more extensively a model is pretrained, the less additional benefit bootstrapping provides. Our findings provide practical insights for efficient language model training and raise important considerations for the reuse of overtrained models.
1 Introduction
Large language models (LLMs) have recently shown astounding performance in various natural language processing tasks, reaching human-level capabilities in some cases achiam2023gpt ; claude3 ; geminiteam2023gemini . However, training/pretraining these models from scratch is computationally expensive and time-consuming, requiring weeks to months even with powerful GPU clusters. To address this challenge within the pretraining stage itself, researchers have explored strategies for bootstrapping/reusing existing pretrained (base) models for various purposes. These include strategies to learn with new pretraining data, or to increase the model size, without starting from scratch. We refer such strategies that pretrain a model in multiple stages collectively as bootstrapped pretraining.
More specifically, we are interested in bootstrapped pretraining strategies that (1) improve domain-specific performance through continual pretraining (CPT) of the base model (see, e.g., ibrahim2024simple ); or (2) scale model capacity via model growth techniques, which increase the model size reusing base model parameters to speed up training chen2015net2net . These strategies have been shown to be promising at accelerating performance improvement of LLMs while reducing the computational cost compared to pretraining from scratch.
On the other hand, neural networks could lose plasticity after training for a long time, thereby reducing their ability to learn from new data effectively berariu2021study ; ash2020warm ; sodhani2020toward . Grown models may also similarly suffer from bad initialization, resulting in ineffective utilization of their enlarged capacity. Hence, it is unclear if the second stage of pretraining can be effectively applied to models that are overtrained, or whether the scaling behavior remains consistent across different training stages and model sizes.
In this paper, we seek to understand the scaling behavior of bootstrapped pretraining methods by running a multitude of controlled language modeling experiments. We find that, from the performance (cross-entropy loss) perspective, bootstrapping overtrained base models leads to saturation in scaling during the second-stage pretraining. Specifically, we show that this effect is quantifiable via scaling laws: The scaling exponent (with respect to the number of training tokens in the second stage) decreases as the base model is trained for longer periods of time. The decrease is proportional to the logarithm of the number of training tokens invested to the first stage. See Figure 1.
Figure 1: Bootstrapped pretraining with overtrained base models leads to saturation in scaling behavior. Left: $D_{2}$ has power-law scaling. We show scaling behavior of second-stage training tokens ($D_{2}$) for different values of first-stage tokens ($D_{1}$). Middle: Interaction term explains decreasing exponents. The fitted exponents in the left plots are used to fit Equation 5 as a function of $D_{1}$, and are shown to agree well with the functional form. Right: Scaling factor has power-law scaling w.r.t. $D_{1}$. Top: Continual pretraining (CPT) on code data. Bottom: Model growth by stacking.
This saturation effect underscores a fundamental trade-off in multi-stage pretraining strategies: bootstrapping a more heavily pretrained base model does not necessarily lead to better performance. Our scaling laws enable a quantitative assessment of when bootstrapping is beneficial versus when training from scratch is preferable, offering practical guidance for efficient language model training.
Contributions. More technically, our main contributions are summarized as follows:
  • We conduct extensive experiments on language models of various sizes and pretraining tokens/datasets, to study the scaling behaviors of bootstrapped pretraining methods, including continual pretraining and model growth.
  • We formulate and empirically validate possible functional forms to quantify the the scaling behavior of these methods. We find that for a wide range of configuration of these methods, the following scaling relation holds:
    $L(D_{1},D_{2})=AD_{1}^{-\alpha_{1}}D_{2}^{-\alpha_{2}+\alpha_{3}\log D_{1}}+E$
    where $L$ is the validation loss after the second stage, $D_{1}$ and $D_{2}$ are the number of training tokens in the first and second stages, respectively, and $A,\alpha_{1},\alpha_{2},\alpha_{3},E$ are positive constants. We denote the term $\alpha_{3}\log D_{1}$ as the interaction term. The scaling exponent with respect to $D_{2}$ then quantifies the saturation effect of bootstrapped pretraining.
  • We also formulate a joint scaling law incorporating model size in addition to dataset sizes, see Equation 7. Using these results, we discuss the practical implications of the scaling laws for efficient bootstrapping of language models. Particularly, we show how the scaling laws provide guidance on when to train from scratch instead of bootstrapping, and how compute optimality is affected by bootstrapped pretraining.
Notations used throughout the paper are summarized in Appendix A.
1.1 Related Work
Continual pretraining. CPT has been widely used to adapt existing LLMs to specific domains sun2020understanding ; jang2022temporal ; jang2022continual , including code yadav2023continualcode ; zan2022cert ; guo2024deepseek and mathematics gong2022continualpretraining ; shao2024deepseekmath domains, to be studied in this work. Systematic studies at scale are relatively fewer, and our methodology mainly follows ibrahim2024simple .
Model growth. While chen2015net2net was the first model growth work in the deep learning era, the idea of growing neural networks from smaller ones can be traced to the 90s NIPS1989_69adc1e1 ; Fahlman1990TheRC . More recent language model-related model growth methods include gong2019efficient ; chen2021bert2bert ; wang2023learning ; wang2023lemon ; shen2022staged ; evci2022gradmax ; yao2024masked , which were subsequently systematically analyzed in du2024stacking , motivating our choice of model growth methods.
Scaling laws. We focus on modeling only the final loss, which has a power-law like behavior, a generic phenomenon not only occuring in neural networks hestness2017deep ; kaplan2020scaling , but is also observed in other natural as well as man-made phenomena clauset2009power . We shall additionally note that recent scaling law studies attempted to fit the whole loss curve albeit with more sophisticated functional forms tissue2024scaling ; wang2025learning ; qiu2025scaling .
Similar to CPT, the reduced capability of transfer learning of code data from models pretrained on natural language data was observed in hernandez2021scaling , where the authors denoted as ossification. A similar phenomenon was also observed in the pretraining-instruction-tuning pipeline springer2025overtrained . These studies however did not quantify the saturated scaling behavior. Previous fine-tuning scaling studies mikami2022scaling ; zhang2024scaling ; bethune2025scaling did not observe the subtler saturation effects as well, which is perhaps due to the smaller scale (in terms of training tokens) of their experiments.
2 Experimental Setup
Model configuration. We consider decoder-only transformers vaswani2017attention pretrained with an autoregressive language modeling objective. Our architecture is LLaMA-like touvron2023llama , incorporating refinements such as SwiGLU activation functions shazeer2020glu and rotary position embeddings su2024roformer . We use the LLaMA tokenizer with a vocabulary size of 32,000. We consider a suite of model sizes up to 1.1B in our experiments. Table 3 of Appendix B contains the model configuration used in this paper.
Training configuration. All experiments are conducted using the Megatron-LM library shoeybi2019megatron . We train models in mixed precision (bfloat16) using the AdamW optimizer loshchilov2017fixing , with maximum learning rates tuned separately for each model size. The learning rate follows the warmup-stable-decay (WSD) schedule bi2024deepseek ; hu2024minicpm , with the final learning rate decayed to one tenth of the peak value. This schedule enables long training runs and facilitates checkpoint reuse for emulating shorter effective training budgets, reducing computational cost. In Appendix C, we also provide an ablation study with the cosine learning rate schedule to show that they achieve similar performance. We train each model for a number of tokens for up to roughly 200 times the model size in tokens in total (up to a few hundred B tokens overall). Tables 4 and 5 in Appendix summarize the main training configurations. See also Appendix B for further experimental details and Appendix C for training details and evaluation on downstream tasks.
Two stages of pretraining.
In the context of bootstrapped pretraining, we are primarily interested in how the loss behaves with respect to the number of training tokens used in the two stages of pretraining:
  • $D_{1}$: the number of tokens used in the first-stage pretraining (base model).
  • $D_{2}$: the number of tokens used in the second-stage pretraining (CPT or model growth).
Base model.
For the first-stage pretraining, we use the CommonCrawl portion of the Slimpajama-DC dataset shen2023slimpajama , containing 368B tokens in total. The models trained in this stage serve as base checkpoints for the second-stage bootstrapped pretraining experiments.
Continual Pretraining.
CPT refers to further training a pretrained model on a large dataset, typically billions of tokens, with the goal of improving performance on a new domain. We distinguish CPT from instruction tuning, which typically trains with smaller datasets (millions of tokens or fewer), and is thus harder to analyze the scaling behavior. For our experiments, we perform CPT on the base model on two domain-specific datasets: Code corpus: Stack/StarCoder li2023starcoder ; and mathematics corpus: OpenWebMath paster2023openwebmath . Unless otherwise noted, we adopt the same optimizer and learning rate configurations as the first-stage pretraining.
Model growth.
Model growth refers to increasing model capacity by adding new layers and/or expanding hidden dimensions, thereby increasing the number of trainable parameters. This strategy aims to leverage the representations learned during first-stage training, allowing the larger model to accelerate learning in the second stage. We investigate two model growth techniques shown to be the most effective (in terms of adding new parameters in the width and depth directions) in prior work du2024stacking (see also Appendix C.5 for more details):
  • Width expansion: Add new neurons to each layer while using function-preserving initialization (FPI), ensuring that the expanded model initially reproduces the behavior of the smaller model chen2021bert2bert ; du2024stacking .
  • Depth-wise stacking: Add new layers to the top of the transformer by copying existing layer weights, thereby extending model depth chen2021bert2bert ; du2024stacking .
After performing this procedure, the larger models are trained on the same dataset as the base model. For both bootstrapped pretraining scenarios in consideration, the validation loss is evaluated on a held-out set from the second-stage dataset. We illustrate the whole framework in Figure 2.
Figure 2: Illustration of bootstrapped pretraining in consideration. Bootstrapped pretraining consists of two stages: (1) first-stage pretraining of a base model for $D_{1}$ tokens on internet/generic data; (2) bootstrapping/second-stage pretraining via continual pretraining or model growth for $D_{2}$ tokens. Section 3 and 4 study and develop scaling laws as a function of these two variables (and additionally model size, NN in Section 5) to predict the final loss after the second stage.
3 Formulating the Data Scaling Laws
In this Section, we derive possible functional forms for the scaling laws of bootstrapped pretraining methods, focusing on the number of training tokens as the main variable of interest. These are then fitted and compared empirically in later sections.
Power-law ansatz.
We begin by assuming that the validation loss $L$ follows the widely observed power-law scaling with respect to a single variable of interest, such as the number of training tokens or the model size hestness2017deep ; hestness2019beyond ; henighan2020scaling :
$L=AX^{-\alpha}+E,$
where $X$ is the variable of interest, $\alpha$ is the scaling exponent, $A$ is the scaling factor, and $E$ is the irreducible loss due to the inherent entropy of the data distribution.  [^1] Strictly speaking, we assume the form $L=\tfrac{A}{(X+1)^{\alpha}}+E$, which ensures finiteness at $X\to 0$. However, since in practice $X\gg 1$ (typically $10^{6}$ or more), we approximate it as $L\approx AX^{-\alpha}+E$ for notational simplicity. Moreover, we use the same symbols $A$, $\alpha$, and $E$ to denote the scaling factor, exponent, and additive constant, respectively, for notational conveniences, although they may differ across different variables of interest.
Our goal is to find a functional form $L(D_{1},D_{2})$ that captures how the second-stage loss depends jointly on both stages. To derive a principled formulation, we impose two natural constraints:
Condition 1: For a fixed base model trained on $D_{1}$ tokens, the second-stage loss should follow a power law with respect to the number of tokens $D_{2}$:
$L(D_{1},D_{2})=L_{D_{1}}(D_{2})=AD_{2}^{-\alpha}+E.$ (1)
This is consistent with classical neural scaling laws and reflects the expectation that, when starting from a fixed initialization, additional tokens improve performance predictably.
Condition 2: For any fixed value of $D_{2}$, the loss should exhibit power-law behavior with respect to the number of first-stage tokens $D_{1}$: $L(D_{1},D_{2})=L_{D_{2}}(D_{1})=AD_{1}^{-\alpha}+E$. This is consistent with the power-law ansatz and captures the intuition that a better-trained base model (i.e., larger $D_{1}$) should result in lower loss. Moreover, this condition implies that as $D_{2}\to 0$, the loss should continuously approach that of the base model: $\lim_{D_{2}\to 0}L(D_{1},D_{2})=AD_{1}^{-\alpha}+E$, which is a natural requirement for:  [^2] As in Footnote 1, we omit the +1+1 term in $D+1$ expressions for notational simplicity, such that terms proportional to $D_{2}$ are finite as $D_{2}\to 0$.
  • CPT: The second stage’s initial loss should begin from the base model’s (evaluated on second-stage dataset).
  • Function-preserving model growth: When model capacity is expanded but initialized carefully, the initial loss should remain close to the base model’s loss.
Together, these conditions lead to the following candidate formulations that jointly satisfy both requirements:
$\textbf{Multiplicative:}\quad L(D_{1},D_{2})$ $=AD_{1}^{-\alpha_{1}}D_{2}^{-\alpha_{2}+\alpha_{3}\log D_{1}}+E.$ (2)
$\textbf{Additive:}\quad L(D_{1},D_{2})$ $=AD_{1}^{-\alpha_{1}}+FD_{2}^{-\alpha_{2}}+E.$ (3)
$\textbf{Hybrid:}\quad L(D_{1},D_{2})$ $=\big(AD_{1}^{-\alpha_{1}}+F\big)D_{2}^{-\alpha_{2}}+E.$ (4)
Note that since $D_{1}^{-\alpha_{1}}D_{2}^{-\alpha_{2}+\alpha_{3}\log D_{1}}=D_{1}^{-\alpha_{2}+\alpha_{3}\log D_{2}}D_{2}^{-\alpha_{2}}$, the multiplicative form with an interaction term satisfies our conditions, while other forms do not. We further note that multiplicative scaling laws have been observed in mikami2022scaling ; zhang2024scaling but without the interaction term. liewscaling empirically showed that sparse upcycling (training sparse mixture-of-experts (MoE) models reusing existing dense models) follows a scaling law similar to Equation 2 but under different motivation and conditions. Our work directly extends their findings to other pretraining paradigms. Equation 4 has also been studied in barnett2024empirical .
4 Data Scaling Laws
Fitting the scaling laws. To determine the functional form described in the previous Section that best describes the scaling behavior of bootstrapped pretraining methods, we use a model of size 0.1B, trained on a $5\times 5$ grid of $D_{1},D_{2}$. For CPT, we consider both code and mathematics datasets, while for model growth, we consider both width expansion and depth-wise stacking that double the size of the base model.  [^3] We increase the size of the hidden dimension by $\sqrt{2}$ for width expansion; and we double the number of non-embedding layers for stacking.
Running the experiments as described above and obtaining the results, we fit the losses with the functional forms of Equations 2, 3, and 4. Following hoffmann2022training ; besiroglu2024chinchilla , we perform optimization using the Huber loss ($\delta=10^{-3}$) and the BFGS algorithm, to fit the logarithm of the loss via the LogSumExp trick applied to the RHS of functional forms. The leave-one-out root mean square error (RMS) serves as the goodness-of-fit metric.
As can be seen in Table 1, the multiplicative form with interaction, Equation 2, consistently achieves the lowest error across all methods and datasets, indicating that it best captures the underlying scaling behavior.
Bottom-up empirical evidence. To strengthen our proposal for Equation 2, we demonstrate through a bottom-up observational approach (instead of the top-down approach of deriving the scaling laws by imposing suitable conditions) that the multiplicative nature of the scaling law, including the interaction term, arises naturally from empirical data.
In Figure 1, we first perform power-law fits to show that the loss follows Equation 1 for fixed $D_{1}$, validating Condition 1 we impose for deriving the scaling laws. More importantly, we observe that the fitted scaling exponent of Equation 1 decreases as $D_{1}$ increases, with the model growth method more pronouncedly so. To quantify this trend, we express (the minus of) the scaling exponent as a function of $D_{1}$, denoted by $\alpha(D_{1})$. A scatter plot of $\alpha(D_{1})$ reveals a clear logarithmic dependence:
$\alpha(D_{1})=\gamma\log D_{1}+E^{\prime}.$ (5)
This relationship fits the empirical data well, as shown in the same Figure. Substituting Equation 5 into Equation 1, we arrive at a term of the form $D_{2}^{-E^{\prime}+\gamma\log D_{1}}$, providing direct empirical justification for the interaction term. This demonstrates that the interaction term is not a theoretical artifact, but rather a necessary component for capturing observed trends in the data. Additionally in the same Figure, we show that the multiplicative dependence, $A\propto D_{1}^{-\alpha_{1}}$ also arises from empirical observations.
We further show that the loss follows a power law with respect to $D_{1}$ for fixed $D_{2}$ in Figure 3 (with more plots in Appendix D). This validates Condition 2 we impose for deriving the scaling laws.
Figure 3: $D_{1}$ has power-law scaling. We show scaling behavior of first-stage training tokens ($D_{1}$) for different values of second-stage tokens ($D_{2}$), indicating that $D_{1}$ also has power-law scaling. Left: Continual pretraining on code data; Right: model growth by stacking. More plots in Appendix D.
Table 1: Multiplicative scaling law with interaction consistently achieves lowest error. Leave-one-out RMS error ($\times 10^{-3}$) for fitting the loss for reusing a 0.1B model. Functional forms are from Equations 2, 3, and specific cases with $\alpha_{3}=0$.
RMS $(\times 10^{-3})$ Mul. Mul. ($\alpha_{3}=0$) Add. Hyb.
CPT (code) 1.573 2.095 3.913 2.245
CPT (math) 1.737 3.147 7.443 3.340
Expand 2.790 4.837 9.855 4.841
Stack 2.845 7.818 10.323 7.748
4.1 Interpretations
Several quantitative and qualitative insights can be made from the scaling law. Fixing $D_{1}$, the effective scaling factor for $D_{2}$ becomes $AD_{1}^{-\alpha_{1}}$. As $D_{1}$ increases, this factor decreases, resulting in a lower initial loss at the start of second-stage pretraining, agreeing with the conventional wisdom that better-pretrained base models (larger $D_{1}$) provide stronger initialization for second-stage pretraining.
However, also at fixed $D_{1}$, the effective scaling exponent with respect to $D_{2}$ is given by $\alpha_{2}-\alpha_{3}\log D_{1}$. This implies that as the base model becomes more overtrained (i.e., larger $D_{1}$), the improvement in loss from additional second-stage tokens becomes increasingly marginal. In other words, the returns from increasing $D_{2}$ diminish, manifesting as saturation effects at higher $D_{1}$.
As shown in Table 7 in Appendix D, the fitted values of $\alpha_{3}$ are consistently positive across all settings and datasets, reinforcing the generality of this saturation phenomenon (see also Section 5.1 for additional evidence). Also note that $\alpha_{3}$ is typically at least an order of magnitude smaller than other exponents (with its effects accumulating logarithmically with $D_{1}$), which may explain why it has not been observed widely in previous scaling studies.
5 A Closer Look At The Scaling Behavior
We first perform a more detailed analysis of the scaling behavior of bootstrapped pretraining methods, by examining more variants of scaling law and training methods. The aim is to validate the robustness of the multiplicative scaling law with interaction. We then propose and validate a joint scaling law of dataset and model sizes. The rest of the Section is subsequently dedicated to studying the practical implications of the scaling laws.
5.1 Variants of Bootstrapped Pretraining
Variant of CPT scaling law. When the dataset used for CPT is the same as the base model, we expect the scaling formula shown below holds:
$L(D_{1},D_{2})=A(D_{1}+D_{2})^{-\alpha}+E$ (6)
Hence, one may expect this holds for CPT on a different dataset as well, as assumed in que2024d ; wang2025learning . Fitting the above formula for the code (math) CPT scenario, we find that the RMS error is 0.0213 (0.0235), higher than the best ones in Table 1. Hence, our scaling law models CPT better, in contrast to previous assumptions, which may have overlooked the overtraining effects on the model.
Variants of CPT. We also consider CPT variants that are commonly used in practice ibrahim2024simple :
  • CPT with replay: A portion of the first-stage data is mixed into the second-stage data, which is common for mitigating catastrophic forgetting. We consider a replay ratio of 0.25, i.e., 25% of the second-stage data is from the first stage.
  • CPT from stable phase: We consider CPT starting from a base model checkpoint in the stable phase of the WSD learning rate schedule. This is to avoid adverse effects from re-warming the learning rate from a decayed value.
Variants of model growth. We further consider model growth variants.
  • Larger growth factor: We consider a larger growth factor of 4 times (instead of 2).
  • Model growth from stable phase: Similar to CPT, we train the stacked model starting from a base model checkpoint in the stable phase of the WSD learning rate schedule. This also avoids re-warming the learning rate from a decayed value.
We again conduct experiments using a 0.1B model, and fit the scaling laws with all possible functional forms as in Section 4 for the above variants. Results are summarized in Table 2, with fitted scaling exponents also shown in Table 7 of Appendix D. These results show that the propsed scaling law still has the lowest RMS, and variants like replay and larger growth factor do not change the functional form. Furthermore, the saturation effects are unlikely to be artifacts of re-warmings of the learning rate. Overall, the proposed scaling law is robust across datasets, configurations, and methods investigated.
Table 2: Multiplicative scaling law with interaction is robust over various configurations (RMS Error). Leave-one-out RMS error ($\times 10^{-3}$) for fitting the loss for variants of bootstrapped pretraining. See Table 1 for definitions and Section 5 for variant details.
RMS $(\times 10^{-3})$ Mul. Mul. ($\alpha_{3}=0$) Add. Hyb.
CPT (replay) 0.987 2.222 4.646 2.223
CPT (stable) 1.371 2.366 4.315 2.367
Stack (x4) 2.152 6.557 8.866 6.667
Stack (stable) 2.957 8.757 10.559 8.822
5.2 Joint Scaling Incorporating Model Size
Let us first focus on CPT where the model size remains unchanged after bootstrapping. We proceed to extend the data scaling law to include model size $N$. To determine the functional form, we impose that the loss follows the well-grounded “Chinchilla" scaling law hoffmann2022training (which jointly models the base model’s loss with respect to dataset and model sizes) with respect to $D_{2}$ and $N$:
$L_{D_{1}}(D_{2},N)=AD_{2}^{-\alpha}+N^{-\beta}+E.$
The straightforward functional form that satisfies the Chinchilla-style scaling law is:
$L(D_{1},D_{2},N)=AD_{1}^{-\alpha_{1}}D_{2}^{-\alpha_{2}+\alpha_{3}\log D_{1}}+BN^{-\beta}+E,$ (7)
which is also consistent with the conditions (power-law ansatz) imposed in Section 3.
Fitting the joint scaling law. We conduct experiments with base models of sizes 15 million (M), 44M, 0.1B, 0.2B, 0.5B and 1B, training them with different numbers of first and second-stage tokens as in Section 4. To fit the joint scaling law, we use the same fitting procedure as in Section 4, additionally taking $N$ as an additional variable. In the left panel of Figure 4, we show that the formula fits the data well for CPT on code, and the goodness-of-fit to the validation data points also shows that it can be extrapolated to larger model and dataset sizes. See Table 8 in Appendix D.2 for the fitted coefficients. In the same Appendix, we further extend our results to model growth where model sizes increase.
Figure 4: Left: Joint scaling law fit for continual pretraining on code data. Orange points indicate the 10% of data with lowest losses used for validation. Right: Model growth efficiency decreases with sunk cost and model size. For token budgets above the curve(s), training from scratch is more efficient than stacking-based model growth; for budgets below, model growth remains advantageous. Shown are the numerical (blue) and analytical (red) solutions of Equation 8.
5.3 Practical Implications
We consider two practical implications based on the extrapolation of the scaling laws.
Training from scratch versus bootstrapped pretraining. We show that by comparing the scaling laws with the base model scaling laws, one can determine whether bootstrapped pretraining is more efficient than training from scratch under different circumstances.
A key practical consideration is how much of the initial investment in pretraining, the so-called sunk cost ($D_{1}$), can be effectively leveraged in the second stage of training. This question has been explored in the context of MoE upcycling strategies komatsuzaki2022sparse ; liewscaling . For example, it was found that upcycling yields benefits up to 120% of the sunk cost. That is, to match the performance of an upcycled MoE model that underwent an additional 0.4 trillion tokens of training after an initial 2T tokens, training a comparable MoE from scratch would require 2.4T tokens—representing an effective saving of 2T tokens.
To investigate this in model growth, we define $D^{*}$ as the number of tokens required for training from scratch to match the performance of a grown model with the same sunk cost, following liewscaling :
$L^{\rm scratch}_{2N}(D^{*})=L^{\rm grown}_{N}(D_{1}=D^{*},D_{2}=D^{*})$ (8)
where $L^{\rm scratch}_{2N}(D)$ is the loss of a model of size $2N$ trained from scratch for $D$ tokens, and $L^{\rm grown}_{N}(D_{1},D_{2})$ is the loss of a model of size $2N$ grown from a base model of size $N$ trained for $D_{1}$ tokens, and then trained for $D_{2}$ tokens. To obtain $L^{\rm scratch}_{2N}(D)$, we fit the losses of models trained from scratch with the Chinchilla-style scaling law (see Table 8 in Appendix D.2 for the fitted coefficients).
The right panel of Figure 4 shows that $D^{*}$ decreases with increasing model size, with $D^{*}$ equal to 13T tokens for a 100B model. When $D_{2}\lesssim D^{*}$, the required from-scratch tokens to catch up is more than 100% of $D_{1}$: model growth remains more efficient than training from scratch. However, beyond this threshold, from-scratch training becomes more efficient. Solving Equation 8, we can approximate the threshold $D^{*}$ analytically as $D^{*}\simeq 13\left(\frac{N}{10^{11}}\right)^{-0.6-0.04\log(N/10^{11})}\;\text{T tokens}$.
For CPT on code data, we can perform similar extrapolation as well, finding that a 7B model trained from scratch begins to outperform bootstrapped models (pretrained on 10T tokens) once more than 300B tokens of domain-specific data are used. See Appendix E for details. These findings offer quantitative guidance for deciding between bootstrapped pretraining and training from scratch (or from intermediate checkpoints), depending on the extent to which the base model is overtrained, as well as the available budget.
Compute optimality of bootstrapped pretraining. For scaling studies of LLMs, the cost of training is often estimated using floating point operations (FLOPs), which we simply refer to as compute. The compute cost can be approximated as $C=6ND$ kaplan2020scaling . One is often interested in training LLMs with the least amount of compute. For CPT, optimizing the compute based on Equation 2 leads to the scaling relations:
$D^{\rm opt}_{2}$ $\propto C_{2}^{\frac{\beta}{\beta+\alpha_{\rm eff}}},\;\;N^{\rm opt}\propto C_{2}^{\frac{\alpha_{\rm eff}}{\beta+\alpha_{\rm eff}}}$ (9)
where $\alpha_{\rm eff}\coloneq\alpha_{2}-\alpha_{3}\log D_{1}$. Notably, as $D_{1}$ increases, $\alpha_{\rm eff}$ decreases, meaning larger models require more tokens for compute-optimal bootstrapped pretraining. We give a similar derivation for model growth in Appendix E. We leave empirical verification of these relationships to future work, primarily due to cost considerations (see Appendix B.4 for an estimate of GPU hours spent in the work).
6 Discussion and Conclusion
In this paper, we have shown that while bootstrapped pretraining methods can accelerate learning, they exhibit scaling saturation as the base model is overtrained. Our work bears practical implications as even relatively tiny models are overtrained to trillions of tokens nowadays zhang2024tinyllama ; dubey2024llama , and bootstrapped pretraining is widely used in practice. We encourage the community to report the amount of pretraining data used for the base model, and even make intermediate checkpoints available, which can be more useful than the final model for bootstrapped pretraining.
We have provided a fairly broad scaling study of two-stage pretraining and there are many potential future directions worth diving deeper into. Let us highlight a few of them.
Incorporating other factors into the scaling laws. We have not considered the scaling behavior with respect to other method-specific factors, such as the CPT replay ratio que2024d ; wang2025learning and the growth factor in model growth du2024stacking , as (1) our focus is on the more broadly applicable scaling behavior (i.e., scaling with respect to training tokens), and (2) adding more factors would increase the complexity of the scaling laws and the number of parameters to fit (and hence the amount of additional experiments/GPU hours/costs needed). We nevertheless expect that these factors can be incorporated quite naturally into our scaling laws which are power-law based.
Theory. Providing a theoretical explanation of the scaling laws is an important open question in the field, but existing work has been focusing on relatively simple theoretical setups like power-law random feature models to reproduce power-law scaling in deep neural networks maloney2022solvable ; bordelon2024dynamical ; paquette20244+ . To our best knowledge, there is no existing theory that can explain the change of scaling exponents with respect to the amount of pretraining data. In the current work, we have focused on the practical aspects of the scaling laws, following previous influential scaling law work kaplan2020scaling ; hoffmann2022training , and have not attempted to provide a theoretical explanation of the scaling laws as the current theoretical understanding of scaling laws is still limited. Nevertheless, it would be interesting to study if such a change of scaling exponents can be explained by existing theories or if new theoretical frameworks are needed.
Appendix A Notations
We summarize main notations used in the paper.
  • $L$: Cross-entropy loss or the (natural) logarithmic loss
  • $D$: Dataset size in token
  • NN: Non-embedding model size or number of non-embedding parameters
  • $A,B,F$: Scaling factors of the power law, independent of the variable under consideration
  • $E$: Irreducible loss of the power law, independent of the variable under consideration
  • $\alpha,\beta,\gamma$: Scaling exponents of the power law, independent of the variable under consideration
  • $n_{\rm layer}$: number of layers of a model
  • $d_{\rm model}$: hidden dimension size of a model
  • $d_{\rm MLP}$: intermediate hidden dimension size of a model
Appendix B More on Architecture and Experimental Design
B.1 Megatron-LM Configuration
Infrastructure. Our experiments are performed on multiple nodes, each consisting of 8 NVIDIA H100 80 GB GPUs, interconnected via InfiniBand HDR. The software we use for training is the Megatron-LM library [57].
We use and modify the Megatron-LM (core v0.8.0) library for our experiments [^4] https://github.com/NVIDIA/Megatron-LM. Models are trained with data type bfloat16. Other optimization libraries used include FlashAttention [16] and TransformerEngine [^5] https://github.com/NVIDIA/TransformerEngine. See the example scripts provided in supplementary material.
B.2 Model Configuration
Let us elaborate more on our model configuration. The intermediate hidden dimension size, $d_{\rm MLP}$, is set to be four times the hidden dimension size, i.e., $4d_{\rm model}$. Bias is not used in the linear layers. We do not consider efficiency-motivated implementations like grouped query attention as well. Attention head number is chosen to increase with model size following standard practices. Other designs of the architecture follow Llama2’s closely [63].
We vary model sizes keeping the ratio $n_{\rm layer}/d_{\rm model}$ to lie in the range 32 to 64, as in [37]. Smaller models for used for ablation studies. See Table 3 for the exact numbers for the model configuration.
Table 3: Models used in our study and their parametric details. Note that $d_{\rm MLP}$, is set to be $4d_{\rm model}$.
Model $n_{\rm layer}$ $d_{\rm model}$ $n_{\rm head}$ NN
15M 9 320 4 14,751,680
44M 12 480 8 44,248,800
0.1B 15 640 8 98,323,840
0.2B 21 832 8 232,623,040
0.5B 26 1,120 16 521,889,760
1B 30 1,504 16 1,085,859,424
B.3 Training Configuration
As discussed in the main text, we adopt the WSD learning rate schedule for all experiments. The number of warmup steps is set approximately equal to the model size, as suggested by [49]. In the final phase of training, the learning rate decays linearly to 10% of its peak value, with the decay phase spanning roughly 10% of the total training steps, following the setup in [27]. To emulate varying token budgets, we save intermediate checkpoints at logarithmically spaced intervals.
The general training configuration is summarized in Table 4, while model-specific hyperparameters, such as warmup iterations, initialization standard deviation ($\sqrt{2/5d_{\rm model}}$ [39]), maximum training steps, batch size, and tuned learning rate, are provided in Table 5. Batch size is scaled with model size according to standard practice, without tuning for optimality. The number of tokens is increased with model size as well, such that the parameter-to-token ratio remains roughly constant across different model sizes. 6 We note that no specialized techniques for mitigating training instabilities, such as Z-loss or QK normalization, are employed, as such instabilities do not arise in our experiments.
Table 4: Training configuration used throughout the paper.
Configuration Details
Context length 1,024
Embedding tying False
Optimizer AdamW [43]
Adam $\beta_{1}$ 0.9
Adam $\beta_{2}$ 0.95
Adam $\epsilon$ 1e-8
Weight decay 0.1
Gradient clipping 1.0
Table 5: Model-dependent training configuration. "init. size" refers to the standard deviation of the normal distribution used for initializing the weights. "Max iter." refers to the maximum iteration run on the model.
Model warmup iter. init. size Max iter. batch size LR
15M 200 0.035 17,600 128 8e-3
44M 200 0.029 17,600 256 4e-3
0.1B 200 0.025 17,600 512 4e-3
0.2B 400 0.022 35,200 512 2e-3
0.5B 800 0.019 70,400 512 4e-4
1B 800 0.016 70,400 1024 4e-4
B.4 GPU Hours and Costs
Instead of reporting the actual runtimes on our cluster, which varied in our experiments due to many factors affecting the cluster (number of available nodes, congestion, etc.), we give a theoretical estimate of total GPU hours used for obtaining the joint scaling law, which involves running the largest tested model with most training tokens in this paper.
The estimate is as follows. We calculate the FLOPs for training the largest first and second-stage models with maximum iterations using the $6ND$ approximation, ignoring the additional FLOPs required to continued pretrain models with shorter iterations (as we can reuse the intermediate checkpoints). We further assume that the per-second TFLOPs of the GPU is 400. We obtain around $10^{9}$ TFLOPs. Taking into account of additional experimentation and ablation runs, we estimate that around 3000 GPU hours were used for the entire project. Assuming a cost of 2 USD per GPU hour, the total cost is around 6000 USD.
Appendix C Training Details
C.1 Ablation of Learning Rate Schedules
Here, we compare the performances of using WSD and the commonly used learning rate (LR) cosine schedules (decaying to 10% of the peak LR value) [63]. The model size we use is 0.1B, with the following training configuration: batch size 512, 4,000 training iterations, and 200 warmup iterations. We can see from Figure 5 that both schedules yield similar (with WSD achieving slightly better final loss) performances. This justifies our choice of using the WSD schedule throughout the paper.
Figure 5: WSD vs cosine LR schedule. We show that the WSD LR schedule achieves similar (even slightly better) final loss compared to the more commonly used cosine LR schedule.
C.2 Ablation of Data Repetition
During model growth, we have used the same data as the base model for the second stage of training. This is because the base model is often trained with a large amount of data, and it is not always possible to obtain a large amount of additional data. However, this means that the second stage of training involves repeating the same data. To study the effect of data repetition, we have made an ablation study, where we compare the performance of using the same data as the base model and using different portion of the Slimpajama data. As shown in Figure 6, we find that the difference is small.
Figure 6: Ablation of data repetition during model growth. We compare the performance of using the same data as the base model and using different data for the second stage of training. The difference is small, indicating that data repetition does not affect the scaling behavior significantly.
C.3 Evaluation with Standard Benchmarks
We compare the performance of our trained 1B base model against existing models with similar sizes, Pythia [9] and TinyLlama [73], based on standard natural language processing benchmarks, ARC [14], lambada [46], logiqa [42], piqa [10], sciq [68], and winogrande [52].
Table 6 shows the results. We see that the model perform similarly to the open models, indicating that our models have been trained correctly.
Table 6: Benchmarks’ performance comparison across models. Reported scores are accuracies (normalized by byte length whenever applicable). The first two columns are scores of existing models. The last column is evaluation results of the largest base model trained with the most number of tokens in this work. Our models are evaluated with the LM Evaluation Harness v0.4.0 library [22].
Models Pythia-1B TinyLlama-1.1B Our 1B model
Datasets Pile Slimpajama & Starcoder Slimpajama
Tokens 100B 103B 74B
ARC-c 25.59 24.32 27.65
ARC-e 47.26 44.91 52.10
lambada 53.52 - 45.08
logiqa 29.49 - 26.11
piqa 69.31 67.30 65.89
sciq 77.3 - 78.10
winogrande 51.22 53.28 54.93
Avg. 50.53 - 49.98
C.4 Continual Pretraining Details
Using a lower LR for CPT is also a common practice [34]. As an ablation, we have experimented with using a constant LR schedule setting the value to be the final LR for the first-stage base model training, on a 0.1B model. We find that it yields worse performance (see Figure 7). Hence, we use the same LR (and LR schedule) as the base model for CPT.
C.5 Model Growth Details
For completeness, we provide more details on the model growth methods used in this work.
Width expansion details. Let us elaborate more on the width expansion method used in this work. We first give a more precise defintion of function preservation. Let $F$ be a function and $G$ as the growth operator. The function preservation condition is defined as:
$F(x)=G(F)(x),\forall x\in\mathcal{X}$
where $\mathcal{X}$ is the input space.
When performing width expansion, the neuron values of each layer are expanded by duplicating the weights of existing neurons, and dividing the output weights by the growth factor.
Stacking details. We first note that stacking does not preserve function but is empirically found to work well in practice [17]. We use the recommended stacking procedure in [17], which is, letting $M$ be the non-embedding part of the base model, the stacked model with growth factor $k$ is given by:
$M^{\prime}=M\circ M\circ...\circ M\quad(k\text{ times})$
where $\circ$ denotes function composition. The embedding and final layer are then simply copied from the base model.
Low LR. We also experiment with using a lower constant LR for stacking as above, and find that it also yields worse performance (see Figure 7). Hence, we use the same LR (and LR schedule) as the base model for stacking as well.
Figure 7: Bootstrapped pretraining with lower LR. We show that using the same LR as the base model achieves better final performance than using a lower LR for a 0.1B model continual pretrained on code data (left) and stacking (right).
Appendix D Scaling Laws Details
D.1 Fitted Exponents and Other Results
We show the fitted exponents of the multiplicative scaling laws studied in Table 1 and 2 in Table 7.
We further show that $D_{1}$ has power-law scaling in Figure 8, for CPT on mathematics data and model growth by expansion, which justifies the multiplicative form of the scaling laws.
Figure 8: Left: $D_{1}$ has power-law scaling. We show scaling behavior of fist-stage training tokens ($D_{1}$) for different values of second-stage tokens ($D_{2}$), indicating that $D_{1}$ also has power-law scaling. From left to right: Continual pretraining on mathematics data, and model growth by expansion.
Table 7: Fitted exponents for the multiplicative scaling law. Corresponding $\alpha_{1},\alpha_{2},\alpha_{3}$ values for the multiplicative interaction fits shown in Table 7.
Variant $\alpha_{1}$ $\alpha_{2}$ $\alpha_{3}$
CPT (code) 0.106 0.146 0.004
CPT (math) 0.981 0.388 0.017
Expand 0.549 0.852 0.024
Stack 0.515 0.350 0.017
CPT (replay) 0.424 0.626 0.018
CPT (stable) 0.920 0.156 0.004
Stack (x4) 0.644 0.829 0.028
Stack (stable) 0.891 0.507 0.009
D.2 Joint Scaling Law
For model growth, we also fit the joint scaling law in Equation 7. Note that there is ambiguity in defining the model size $N$ in the context of model growth. We consider $N$ to be the size of the model before growth, and as we keep the growth factor fixed to be 2 in our experiments, the new model size after growth is $N^{\prime}=2N$. Therefore, $N$ and $N^{\prime}$ differ by a constant factor of 2, which can be absorbed into the coefficient $B$ in Equation 7. Unless stated otherwise, we use $N$ in the fitting of the joint scaling law for model growth. In Figure 9, we show the fit of the joint scaling law for stacking.
Figure 9: Fits of the joint upcycling scaling law of stacking. 10% of the collected data points with lowest losses are used for validation (orange points).
We provide the fitted coefficients of the joint scaling law in Table 8. In addition to bootstrapped pretraining, we also show the fitted coefficients for base models trained from scratch on the same dataset as the second stage (code data for CPT, and the same data as the base model for model growth). For model growth, we consider a growth factor of 2, and fit the parameters with model size $N$ before stacking for comparison conveniences. We further note that the fitted coefficients are produced by fitting the joint scaling law to all data points collected, including those used for validation in Figure 9.
Table 8: Fitted coefficients for joint scaling laws. Note that for stacked models, we fit the coefficients with model size $N$ before stacking for comparison conveniences.
$\alpha/\alpha_{1}$ $\alpha_{2}$ $\alpha_{3}$ $\beta$ AA $B$ $E$
Base (Slimpajama) 0.092 - - 0.105 10.383 10.085 0.041
From-scratch (code) 0.113 - - 0.234 8.143 27.286 0.105
CPT (code) 0.048 0.126 0.001 0.238 15.062 27.234 0.105
Stack 0.087 0.119 0.003 0.173 33.394 22.471 0.041
D.3 Why Fitting Scaling Laws Separately?
We justify our decision to fit two separate scaling laws—one for models trained from scratch and another for bootstrapped pretraining—instead of employing a unified formulation that spans all stages.
First, the scaling behavior of models trained from scratch is expected to differ from that of models undergoing bootstrapping. Specifically, at the limit $D_{1}\to 0$, grown models are initialized from a pretrained base model, whereas scratch-trained models start from random weights. This difference in initialization leads to distinct learning dynamics and, consequently, different scaling law parameters.
Similarly, the scaling behavior of the base model differs from that of the second-stage training in the limit $D_{2}\to 0$. As shown in Figure 7, second-stage training often begins with a rewarming phase, during which the loss initially increases before decreasing. This early instability deviates from the expected scaling of dense models, although the overall trend remains correlated—supporting the validity of Condition 1, due to function preservation and empirical observations that rewarming does not entirely disrupt loss behavior (as quoted from [15]: “In practice, few empirical phenomena obey power laws for all values of xx”).
Finally, our preliminary experiments indicate that a unified scaling law does not provide a satisfactory fit across both stages. We therefore opt to model them separately. We also note that there may exist alternative functional forms that could better capture the full range of behavior, but they may violate some of the well-established conditions for scaling laws (power law), and may overcomplicate the analysis or overfit the data; hence, we leave their exploration to future work.
Appendix E More Practical Implications
To further examine the trade-off between training from scratch and bootstrapped pretraining, we fit Chinchilla-style scaling laws to models trained from scratch on the same datasets used in our second-stage experiments, namely, code or math data for CPT, and the original dataset for model growth. In Figure 10 and 11, we plot the validation loss as a function of training tokens for both approaches, across a range of sunk costs, using a fixed model size of 7B. For CPT on code data, we find that models trained from scratch begin to outperform bootstrapped models (pretrained on 10T tokens) once more than 300B tokens of domain-specific data are used.
Figure 10: Tokens required for from-scratch training to catch up with bootstrapped pretraining. We compare loss-versus-token plots of from-scratch and bootstrapped pretraining at various base model training budgets (sunk costs). Stacking is only considered efficient when $D_{\rm scratch}<$ sunk cost. We observe that the efficiency of stacking diminishes with sunk cost.
Figure 11: Tokens required for from-scratch training to catch up with bootstrapped pretraining. Conclusion similar to those made in Figure 10 can be made for CPT on code data, where the scaling efficiency decreases with first-stage training tokens.
E.1 Compute Optimality of Model Growth
For model growth with growth factor 2, we want to scale $N_{2},D_{2}$ optimally, $N^{\rm opt}_{2},D^{\rm opt}_{2}$, given a FLOPs budget and fixing $D_{1}$, while minimizing the loss $L$, which we write as $L_{D_{1}}(D_{2},N_{1})$. Here, $N_{1}$ is the model size before growth, and $N_{2}=2N_{1}$. This is equivalent to solving the following:
$\left.\frac{\partial}{\partial D_{2}}L_{D_{1}}(D_{2},C_{2}/12D_{2})\right|_{D_{2}=D_{2}^{\rm opt}}=0,$
$\left.\frac{\partial}{\partial N_{1}}L_{D_{1}}(C_{2}/12N_{1},N_{1})\right|_{N_{1}=N^{\rm opt}_{1}}=0$
where we have used $N_{2}=2N_{1}$ and $C_{2}=6N_{2}D_{2}$. Solving the above equations leads to
$D^{\rm opt}_{2}$ $=G\left(\frac{C_{2}}{12}\right)^{a},$
$N^{\rm opt}_{1}$ $=G^{-1}\left(\frac{C_{2}}{12}\right)^{b}$
where
$G$ $\coloneq\left(\frac{A_{\rm eff}\alpha_{\rm eff}}{B\beta}\right)^{1/(\alpha_{\rm eff}+\beta)}$
$a$ $\coloneq\frac{\beta}{\alpha_{\rm eff}+\beta}$
$b$ $\coloneq\frac{\alpha_{\rm eff}}{\alpha_{\rm eff}+\beta}$
$A_{\rm eff}$ $\coloneq AD_{1}^{-\alpha_{1}}$
$\alpha_{\rm eff}$ $\coloneq\alpha_{2}-\alpha_{3}\log D_{1}$
We can henceforth relate $D^{\rm opt}_{2}$ and $N^{\rm opt}_{1}$ via
$D^{\rm opt}_{2}=G\left(GN^{\rm opt}_{1}\right)^{a/b}\propto\left(N^{\rm opt}_{1}\right)^{\beta/\alpha_{2}-\alpha_{3}\log D_{1}}$
and
$N^{\rm opt}_{1}=G^{-1}\left(G^{-1}D^{\rm opt}_{2}\right)^{b/a}\propto\left(D^{\rm opt}_{2}\right)^{(\alpha_{2}-\alpha_{3}\log D_{1})/\beta}$